Question 18 0/1 pts Russin Tax Planning Service bought computer equipment for $21,000 on January 1, 2016. It has an estimated why is 3938 the right answer?
Answer #1

Depreciation for the year 2016 = Cost/useful life = 21000/4years = $ 5250

The required depreciation is for the period of January 1,2016 to September,2016. That would be period of 9 months. The proportional depreciation amount needs to be debited for 9 months.

Depreciation for 9 months = 5250 * 9/12 = $ 3938